Find the maximum slope of the line tangent to the curve y = -x³+6x²+16x+20.
Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

I've seen this problem (with different numbers) at least twice earlier today.
Look at the solved problems.

Answer by stanbon(75887)

Find the maximum slope of the line tangent to the curve y = -x³+6x²+16x+20.
----
Slope at every point = -3x^2+12x+16
--------
To find max slope take the derivative and solve::
-6x+12 = 0
x = 2
-------
Ans: Max Slope = -3*2^2+12*2+16 = -12 + 24 + 16 = 28
-----------
Cheers,
Stan H.
-----------